Tickets for a US lottery cost $45 to the overseas participants and $30 to the locals.If the tickets were sold for a total of $12000,how many tickets were sold to the locals? 

Grade:12th pass

3 Answers

Eshan
5 years ago
Dear student,

Let the number of overseas participants bexand locals bey. Hence total cost of tickets sold=45x+30y=12000.

However to obtain the vaules of the two unknown variables , we need one more equation and so one more given condition.
Abubakar Naeem
5 years ago
45x+30(150)=12000
45x=12000-4500
45x=7500
x=7500/45=100 a valid answer not in fraction so the Answer is 150
Tehsina shoaib
3 years ago
45X+30(250)=12,000
45X=12,000-7500
45X=4500
x=100
so correct answer is 250 not in fraction,so the answer is 250